A jeweler examines a diamond with a magnifying glass. If the near-point distance of the jeweler is 20.7 cm, and the focal length of the magnifying glass is 7.00 cm, find the angular magnification when the diamond is held at the focal point of the magnifier. Assume the magnifying glass is directly in front of the jeweler's eyes. (Include the sign.)

A grade-school student plans to build a 55 power telescope as a science fair project. She starts with a magnifying glass with a focal length of 2.8 cm as the eyepiece. What focal length is needed for her objective lens? (Include the sign.)
m

An optical fiber uses one glass clad with another glass. What is the critical angle? (Assume the glass in the fiber has an index of refraction of 1.70, and the glass in the cladding has an index of refraction of 1.55.)

°
Answer #1

1.

Since the jeweler holds the diamond at the focal point of the magnifying glass, its image will be produced infinitely far away. This gives an angular magnification of :

M = ' /

M = N/f = 20.7 / 7

M = 2.96